# Unleashing Large Multimodal Model’s Ability for Painting 3D Objects with Realistic Materials

### v2

@@ 0.D.2 Effects of Different Texture Maps
We validate the impact of various texture maps generated by Make-it-Real on the appearance of 3D objects, as illustrated in Fig. 17. For each example, the first column lacks the corresponding map enhanced by Make-it-Real, while the second column includes the same conditions with the specific material map. Initially, we compare the effect of metalness on object appearance. We observe that the objects on the right, with a higher metalness map value, exhibit higher reflectivity, and the reflected light’s color is similar to the material’s own color, closely resembling real-world objects and appearing more aesthetically pleasing. In contrast, the objects on the left without the metalness map have lower reflectivity, and the reflected light tends to be white.
@@
Appendix 0.E More Qualitive Results
0.E.1 Make-it-Real for Existing 3D Assets
@@
In this section, we show more qualitative results of Make-it-Real for existing 3D assets from [15]. Results are shown in Fig. 18. The first row presents the original 3D object with only a diffuse map, while the second row showcases the object enhanced by our method with various material maps.
0.E.2 Make-it-Real for 3D Generation Models
In this section, we show more qualitative results of Make-it-Real for 3D generation models. We demonstrate the effectiveness of Make-it-Real for refining both image-to-3D and text-to-3D generated objects, as illustrated in Fig. 19 and Fig. 20, respectively.
@@ 0.E.2 Make-it-Real for 3D Generation Models
Figure 18: More qualitative results of Make-it-Real refining existing 3D assets without material. Objects are selected from Objaverse[15] with diffuse only.
Figure 19: More qualitative results of Make-it-Real refining objects generated by TripoSR [50]. Single View images are selected from Objaverse [15] rendered images. First line per two are original generated objects, Second lines are objects refined by Make-it-Real.
Figure 20: More qualitative results of Make-it-Real refining objects generated by Instant3D [30]. Text prompts selected from GPTeval3D [56]. First line per two are original generated objects, Second lines are objects refined by Make-it-Real.
@@
Figure 21: Visualization of generated texture maps. The first column represents the original query diffuse map of 3D objects, while the subsequent columns showcase the corresponding material maps generated by Make-it-Real.
0.E.3 Visualization of generated texture maps.
@@
In this section, we show visualization of generated texture maps in Fig. 21. Our approach, guided by the query diffuse reference, produces material maps with distinct material partitions and maintains a distribution consistent with the diffuse map. As a result, the enhanced object exhibits both realism and texture consistency.

### v3

@@ D.1 Effects of Different Texture Maps
We validate the impact of various texture maps generated by Make-it-Real on the appearance of 3D objects, as illustrated in Figure 16. For each example, the first column lacks the corresponding map enhanced by Make-it-Real, while the second column includes the same conditions with the specific material map. Initially, we compare the effect of metalness on object appearance. We observe that the objects on the right, with a higher metalness map value, exhibit higher reflectivity, and the reflected light’s color is similar to the material’s own color, closely resembling real-world objects and appearing more aesthetically pleasing. In contrast, the objects on the left without the metalness map have lower reflectivity, and the reflected light tends to be white.
@@
Appendix E More Qualitive Results
E.1 Make-it-Real for Existing 3D Assets
@@
In this section, we show more qualitative results of Make-it-Real for existing 3D assets from [15]. Results are shown in Figures 17 and 18. The first row presents the original 3D object with only a albedo map, while the second row showcases the object enhanced by our method with various material maps.
@@
Figure 17: More qualitative results of Make-it-Real refining existing 3D assets without material. Objects are selected from Objaverse[15] with albedo only.
Figure 18: More qualitative results of Make-it-Real refining existing 3D assets without material. Objects are selected from Objaverse[15] with albedo only.
@@
Figure 19: Visualization of generated texture maps. The first column represents the original query albedo map of 3D objects, while the subsequent columns showcase the corresponding material maps generated by Make-it-Real.
Figure 20: Visualization of generated texture maps. We show part of the SVBRDF material maps generated by Make-it-Real and the final rendering results. We displayed the texture maps and the corresponding 3D rendering effects. The albedo is the input, and the following four columns show the material effects in the UV space and on the 3D object.
E.2 Visualization of generated texture maps.
@@
In this section, we show visualization of generated texture maps in Figures 19 and 20. Our approach, guided by the query albedo reference, produces material maps with distinct material partitions and maintains a distribution consistent with the albedo map. As a result, the enhanced object exhibits both realism and texture consistency.
E.3 Limitations and future work.
While promising results have been achieved by MLLMs for texture assignment through Make-it-Real. Our work still faces several unresolved challenges. First, although our method can achieve appropriate texture assignment for albedo-only model, it does not support reverse transform from shaded texture map to albedo map for generated 3D objects. This causes problem of assigning different materials to the dark shadow and highlight area when generated object with mesh already shaded in different lighting conditions. Second, we find base model quality is essential for MLLM to assign correct materials. When base 3D object is of low quality, it is difficult for MLLMs to identify object properties when ground truth text prompt describing the object is not available. Additionally, there has been limited progress in material segmentation in 2D and 3D demoain[48, 36]. The material segmentation algorithm included in our proposed pipeline can serve as a useful baseline and provide inspiration for future work. We also explore the integration of the pipeline with 3D segmentation networks, which shows promising results. For future work, we consider methods mitigating these challenges as well as adding user friendly control into our fully-automatic pipeline.
Appendix F Broader Impacts
Potential positive societal impacts: The proposed method facilitates more realistic and accurate representations of materials in 3D models, benefiting industries such as gaming, virtual reality, and film, leading to more immersive and engaging experiences. By automating the material assignment process, Make-it-Real significantly reduces the time and effort required for 3D content creators, allowing for more efficient workflows and enabling creators to focus on more creative aspects of their work. This approach can democratize high-quality 3D content creation by making advanced material application techniques accessible to a broader range of users, including those without specialized skills in graphic software.
Potential negative societal impacts: The improved realism in 3D assets could be exploited for creating highly convincing fake visuals or deepfakes, which might be used in disinformation campaigns or to mislead audiences. There is a risk that the materials generated could inadvertently perpetuate biases or stereotypes if the training data for GPT-4V includes biased representations of certain materials or objects. As the method involves processing and recognizing visual data, there could be concerns regarding the privacy of any real-world images used as inputs, particularly if they contain sensitive or personal information.